OpenStudy (anonymous):

divide and check write your answer using non negative exponents 25x^3y^2 / 5x^5y^2

OpenStudy (anonymous):

5/x^2

OpenStudy (saifoo.khan):

\[\frac{5}{x^2}\]

OpenStudy (anonymous):

25/5 = 5 y^2/y^2 = 1 x^3/x^5 = 1/x^2 5/x^2

OpenStudy (anonymous):

so when writing the answer i just put 5x^2??

OpenStudy (anonymous):

No, 5/x^2

OpenStudy (anonymous):

for some reason i got 5x^-2

OpenStudy (anonymous):

That is the same thing.

OpenStudy (anonymous):

thank you! :)